# Zhang Q Y, Xu L, Huang J H, et al

### v1

@@ 3.9.1 High-Speed Inter-Satellite Communication
The DSIN requires high-speed inter-satellite communication to achieve global ubiquitous coverage, but the increasing demand for inter-satellite communication bandwidth, coupled with the instability of free-space channels, presents new challenges in DSIN. As the communication capacity required for satellites and various spacecraft grows exponentially, the current communications based on RF are becoming insufficient to meet the sharply rising demand for capacity [265]. Inter-satellite Free Space Optical (FSO) communication has emerged as a promising alternative or complementary solution to traditional RF systems, owing to its vast unlicensed bandwidth and the ability to transmit data at exceptionally high rates over long distances [266]. Moreover, in contrast to conventional RF-based wireless systems, the narrow and directional nature of the laser beams used in FSO communications offers enhanced security, reduced power consumption, and immunity to electromagnetic interference. Constellation systems such as Starlink, OneWeb, Kuiper, and Telesat have already adopted inter-satellite FSO communication as one of their core transmission links. Nevertheless, despite the great potential of FSO communication for DSIN, its performance suffers from various limitations and challenges. On the one hand, the FSO communication is susceptible to the detrimental effects of atmospheric turbulence, including beam-wandering-induced pointing errors, beam scintillation, and attenuation caused by weather conditions such as haze, snow, fog, and clouds [265]. Specifically, pointing errors induced by beam wandering and misalignments between receivers and transmitters can severely impact the reliability of the link. To address these challenges, the hybrid FSO/RF scheme within satellite networks presents a promising solution, leveraging the complementary nature of both technologies to enhance link quality without causing interference between the FSO and RF links [267]. Relay-assisted transmission via high-altitude platforms (HAPs) offers another promising solution to enhance the usability of FSO links. For example, unmanned aerial vehicles (UAVs) can be deployed between satellites and GSs to establish a space-air-ground FSO network within DSIN, providing high reliability and high-speed transmissions. This approach has already been explored in [268] and has been validated by NASA’s Laser Communications Relay Demonstration (LCRD) mission, which successfully transmitted data from GEO at a rate of 1.2 Gb/s. However, the performance of FSO communication significantly deteriorates as the propagation distance increases, particularly in environments with strong turbulence.
@@ 4.2 Resource Management
Looking ahead, the future DSIN will require innovative approaches to enhance resource management efficiency. One promising direction is the development of integrated resource management for terrestrial-satellite systems, which can optimize the use of resources across both domains. Additionally, research into intelligent resource management for satellite and terrestrial spectrum shared networking is essential, focusing on spectrum sensing, prediction, and allocation to improve spectrum efficiency. Moreover, in the distributed satellite network architecture, blockchain-based frameworks for decentralized resource management could enable autonomous resource allocation among CCS system, enhancing flexibility and reducing the need for centralized control. By providing a secure, distributed ledger, blockchain can facilitate transparent and verifiable transactions, improving the reliability and scalability of resource sharing. Further, future DSIN would benefit from the development of adaptive algorithms capable of real-time resource allocation adjustments to handle dynamic demand fluctuations. These algorithms should be designed to operate in high-mobility environments, allowing networks to dynamically adjust to the changing conditions of satellite constellations without compromising efficiency. The last but not the least direction is multi-orbit satellite collaboration, where collaborative resource management strategies among satellites in different orbits, such as LEO, MEO, and GEO can be obtained. This significantly enhances resource utilization by enabling satellites to share resources based on their specific capabilities and operational constraints, resulting in improved network resilience and performance.
@@ 4.4 Testbeds of DSIN
The international application status of these platforms is marked by several key developments and trends. For instance, the integration of Software-Defined Networking (SDN) in satellite networks has been a significant step forward. The SDN allows for the separation of the data plane and control plane, enabling centralized management of network resources and dynamic reconfiguration in response to varying traffic demands and network conditions [347]. This approach not only reduces the computational burden on individual satellites but also optimizes the use of network resources, catering to the high-dynamic and heterogeneous nature of satellite networks. As a cost-effective SatCom solutions for 5G, the virtualisation of SatCom network functions is validated in the SAT5g project [348], ensuring compatibility with the 5G SDN and NFV architecture. To address the challenge of analyzing and evaluating integrated space, aerial, and ground networks, a SAGIN simulation platform which integrates multiple network protocols, node mobility, and control algorithms is developed in [349], optimizing the network functions such as access control and resource orchestration in a combined centralized and decentralized manner. While the flexibility and scalability brought by SDN/NFV technologies are beneficial for heterogeneous network architecture, the issues of deploying the SDN controllers in DSIN and coordinating the actions among the controllers require careful investigation. As a viable option, cloud/edge computing and network slicing can be incorporated into resource isolation, cloud services, computation offloading, and edge caching, fundamentally meet the varying requirements of ever-increasing services and applications. In addition, the development of testbeds like the DVB-RCS2/S2 has been instrumental in the design and validation of next-generation satellite systems [350]. These testbeds provide a distributed environment that mimics the operational characteristics of real satellite networks, allowing researchers and engineers to experiment with various protocols, routing algorithms, and network configurations in a controlled setting.
@@ 5.1 Network Resource Virtualization in DSIN
In NRV-enabled DSIN, heterogeneous resources can be dynamically scheduled and prioritized to address the areas of greatest need, thereby optimizing key performance indicators across multiple satellites. Despite the potential of NRV, several critical challenges hinder its full integration into DSIN, the foremost being the lag effect in resource allocation caused by long-distance communication. The physical distances between satellites introduce latency issues, potentially undermining the effectiveness of NRV, particularly in time-sensitive applications. Research efforts have explored multi-layer architectures and edge computing solutions to reduce latency [359]. By enabling data processing closer to the source, edge computing helps mitigate latency. In satellite networks, LEO satellites can function as edge devices, processing data locally before transmission. However, while edge computing alleviates some latency issues, it only provides a partial solution and demands substantial onboard processing capabilities, posing a challenge for smaller satellites with limited resources. Moreover, achieving resource virtualization across heterogeneous satellite platforms presents challenges in forming a unified, interoperable resource scheduler.
@@
Another critical issue lies in the limitations of multidimensional heterogeneous resources in DSIN, which are confined to different network layers, lacking interoperability and unified global coordination. To address this, the virtualized hybrid satellite-terrestrial systems (VITAL) project introduces the network functions virtualization (NFV) and software-defined networking (SDN) technologies into satellite networks, enabling flexible and resilient resource management in satellite-ground networks [364]. The SDN facilitates dynamic resource management in heterogeneous satellite networks. By dynamically deploying SDN controllers in LEO satellite constellations, it becomes possible to effectively address network traffic fluctuations caused by variations in user geographic locations and time zones [365]. Moreover, the SDN/NFV technologies in SAGIN have been proven to effectively transform the network from a connection-based model to a service-based model, offering on-demand resource allocation and service customisation [366]. However, the inherent resource dynamics and service uncertainties in DSIN render traditional scheduling methods unsuitable for decision-making tasks requiring high efficiency and rapid response, such as service function chain deployment and mapping. Reinforcement learning (RL), as a self-learning and adaptive decision-making approach, has been explored in SDN/NFV networks to some extent [367]. Nevertheless, how to apply RL to integrate lightweight SDN/NFV for enabling adaptive and scalable NRV in DSIN remains relatively scarce, highlighting a gap in the current research.
@@ 5.2 Distributed AI Inference and On-Orbit Information Processing
Given the constrained transmission capacity of satellite-to-ground links, traditional architectures reliant on GS struggle to address the growing demands for massive data return [368, 369]. Moreover, the limited resources of a single satellite render it impractical to process all received data independently. Consequently, leveraging on-orbit computing and storage resources across DSIN becomes essential for the real-time processing of vast sensor data collections. Distributed AI technologies, such as federated learning [370], transfer learning and split learning [371], can be effectively employed within CCS system in DSIN to facilitate on-orbit information processing. Moreover, With the rapid development of AI, large language models (LLMs) have become a star technology in the field of natural language processing. However, the inference process of LLMs often faces significant challenges in terms of computational resources and time costs, especially for on-orbit intelligent applications in satellite networks. To address this issue, distributed AI inference technology has emerged. Distributed AI inference involves breaking down the large language model inference task, which would typically be executed on a single machine, into multiple sub-tasks that can be processed in parallel across multiple nodes with limited computational power. This approach enables the efficient utilisation of computational resources from distributed and edge nodes, significantly improving on-orbit information processing and inference speed, while also reducing inference costs. During the distributed inference offloading process, managing a large number of computational resources, including CPUs, GPUs, and others, is crucial. To accelerate the model inference speed, NRV and distributed satellite computing network architecture, as mentioned in Section $2.2$ and $5.1$, can be employed to achieve compact and efficient computational power collaboration and task priority scheduling. Furthermore, in response to the dynamic network characteristics of DSIN, such as frequent topology changes and narrow communication time windows between the satellites and GSs, distributed incremental inference emerges as a promising technique. By learning the implicit relationships between various tasks or model information, it abstracts a distributed inference model with basic inference capabilities. This approach avoids full-scale learning and model updates, thereby accelerating model inference speed while reducing communication overhead.
@@ 5.3 Semantic Communications in DSIN
The 6G-enabled DSIN are expected to provide ubiquitous intelligent services with stringent QoS requirements in dynamic network environments, which prompts a shift from conventional architectures that focus on high transmission rates between transceivers to new architectures centered on intelligent connectivity of everything. The resulting extensive scales of data which require a massive amount of bits to represent, impose a huge burden and bottleneck on communication systems that handle only bit-level reconstruction, especially on DSIN with limited transmission power, storage space and computing resources [382]. In response, a novel paradigm known as semantic communication is inspired as a potential technology to break the bottleneck by utilizing the semantics and content of data to optimize the utilization of communication resources [383, 384].
@@ 6 Conclusion
This work was supported in part by the Major Key Project of PCL under Grant PCL2024A01, in part by the National Natural Sciences Foundation of China (NSFC) under Grant 62071141 and Grant 62027802, and in part by the National Key R&D Program of China No. 2020YFB1806900, Beijing Natural Science Foundation No. L242012.
@@
3GPP 3rd Generation Partnership Project MEO Medium Earth Orbit
@@
GRAND Guessing Random Additive SDN Software Defined Networking
@@
LLMs Large Language Models

### v2

@@ 3.9.1 High-Speed Inter-Satellite Communication
The DSIN requires high-speed inter-satellite communication to achieve global ubiquitous coverage, but the increasing demand for inter-satellite communication bandwidth, coupled with the instability of free-space channels, presents new challenges in DSIN. As the communication capacity required for satellites and various spacecraft grows exponentially, the current communications based on RF are becoming insufficient to meet the sharply rising demand for capacity [290]. Inter-satellite Free Space Optical (FSO) communication has emerged as a promising alternative or complementary solution to traditional RF systems, owing to its vast unlicensed bandwidth and the ability to transmit data at exceptionally high rates over long distances [291]. Moreover, in contrast to conventional RF-based wireless systems, the narrow and directional nature of the laser beams used in FSO communications offers enhanced security, reduced power consumption, and immunity to electromagnetic interference. Constellation systems such as Starlink, OneWeb, Kuiper, and Telesat have already adopted inter-satellite FSO communication as one of their core transmission links. Nevertheless, despite the great potential of FSO communication for DSIN, its performance suffers from various limitations and challenges. On the one hand, the FSO communication is susceptible to the detrimental effects of atmospheric turbulence, including beam-wandering-induced pointing errors, beam scintillation, and attenuation caused by weather conditions such as haze, snow, fog, and clouds [290]. Specifically, pointing errors induced by beam wandering and misalignments between receivers and transmitters can severely impact the reliability of the link. To address these challenges, the hybrid FSO/RF scheme within satellite networks presents a promising solution, leveraging the complementary nature of both technologies to enhance link quality without causing interference between the FSO and RF links [292]. Relay-assisted transmission via high-altitude platforms (HAPs) offers another promising solution to enhance the usability of FSO links. For example, unmanned aerial vehicles (UAVs) can be deployed between satellites and GSs to establish a structured space-air-ground FSO network within DSIN, providing high reliability and high-speed transmissions . This approach has already been explored in [294] and has been validated by NASA’s Laser Communications Relay Demonstration (LCRD) mission, which successfully transmitted data from GEO at a rate of 1.2 Gb/s. However, the performance of FSO communication significantly deteriorates as the propagation distance increases, particularly in environments with strong turbulence.
@@ 4.2 Resource Management
Looking ahead, the future DSIN will require innovative approaches to enhance resource management efficiency. One promising direction is the development of integrated resource management for terrestrial-satellite systems, which can optimize the use of resources across both domains. Additionally, research into intelligent resource management for satellite and terrestrial spectrum shared networking is essential [362], focusing on spectrum sensing, prediction, and allocation to improve spectrum efficiency. Moreover, in the distributed satellite network architecture, blockchain-based frameworks for decentralized resource management could enable autonomous resource allocation among CCS system, enhancing flexibility and reducing the need for centralized control. By providing a secure, distributed ledger, blockchain can facilitate transparent and verifiable transactions, improving the reliability and scalability of resource sharing. Further, future DSIN would benefit from the development of adaptive algorithms capable of real-time resource allocation adjustments to handle dynamic demand fluctuations. These algorithms should be designed to operate in high-mobility environments, allowing networks to dynamically adjust to the changing conditions of satellite constellations without compromising efficiency. The last but not the least direction is multi-orbit satellite collaboration, where collaborative resource management strategies among satellites in different orbits, such as LEO, MEO, and GEO can be obtained. This significantly enhances resource utilization by enabling satellites to share resources based on their specific capabilities and operational constraints, resulting in improved network resilience and performance.
@@ 4.4 Testbeds of DSIN
The international application status of these platforms is marked by several key developments and trends. For instance, the integration of SDN in satellite networks has been a significant step forward. The SDN allows for the separation of the data plane and control plane, enabling centralized management of network resources and dynamic reconfiguration in response to varying traffic demands and network conditions [374]. This approach not only reduces the computational burden on individual satellites but also optimizes the use of network resources, catering to the high-dynamic and heterogeneous nature of satellite networks. As a cost-effective SatCom solution for 5G, the virtualisation of SatCom network functions is validated in the SAT5g project [375], ensuring compatibility with the 5G SDN and NFV architecture. To address the challenge of analyzing and evaluating integrated space, aerial, and ground networks, a SAGIN simulation platform which integrates multiple network protocols, node mobility, and control algorithms is developed in [376], optimizing the network functions such as access control and resource orchestration in a combined centralized and decentralized manner. While the flexibility and scalability brought by SDN/NFV technologies are beneficial for heterogeneous network architecture, the issues of deploying the SDN controllers in DSIN and coordinating the actions among the controllers require careful investigation. As a viable option, cloud/edge computing and network slicing can be incorporated into resource isolation, cloud services, computation offloading, and edge caching, fundamentally meeting the varying requirements of ever-increasing services and applications. In addition, the development of testbeds like the DVB-RCS2/S2 has been instrumental in the design and validation of next-generation satellite systems [377]. These testbeds provide a distributed environment that mimics the operational characteristics of real satellite networks, allowing researchers and engineers to experiment with various protocols, routing algorithms, and network configurations in a controlled setting.
@@ 5.1 Network Resource Virtualization in DSIN
In NRV-enabled DSIN, heterogeneous resources can be dynamically scheduled and prioritized to address the areas of greatest need, thereby optimizing key performance indicators across multiple satellites. Specifically, based on the distributed satellite computing network architecture presented in Section 2.2, the DSIN can decentralize the computation, storage, and networking resources of satellites across different spatial environments, and connects them via ISL to form a collaborative working framework. This architecture allows for the flexible allocation of satellite, ground, and edge network resources, enabling efficient information collection, processing, computation, storage, and distribution. It is particularly well-suited to meet the massive, intelligent information service demands of the future, which will be driven by the vision of the Internet of Everything.
Despite the potential of NRV, several critical challenges hinder its full integration into DSIN, the foremost being the lag effect in resource allocation caused by long-distance communication. The physical distances between satellites introduce latency issues, potentially undermining the effectiveness of NRV, particularly in time-sensitive applications. Research efforts have explored multi-layer architectures and edge computing solutions to reduce latency [386]. By enabling data processing closer to the source, edge computing helps mitigate latency. In satellite networks, LEO satellites can function as edge devices, processing data locally before transmission. However, while edge computing alleviates some latency issues, it only provides a partial solution and demands substantial onboard processing capabilities, posing a challenge for smaller satellites with limited resources. Moreover, achieving resource virtualization across heterogeneous satellite platforms presents challenges in forming a unified, interoperable resource scheduler.
@@
Another critical issue lies in the limitations of multidimensional heterogeneous resources in DSIN, which are confined to different network layers, lacking interoperability and unified global coordination. To address this, the virtualized hybrid satellite-terrestrial systems (VITAL) project introduces the network functions virtualization (NFV) and SDN technologies into satellite networks, enabling flexible and resilient resource management in satellite-ground networks [391]. The SDN facilitates dynamic resource management in heterogeneous satellite networks. By dynamically deploying SDN controllers in LEO satellite constellations, it becomes possible to effectively address network traffic fluctuations caused by variations in user geographic locations and time zones [392]. Moreover, the SDN/NFV technologies in SAGIN have been proven to effectively transform the network from a connection-based model to a service-based model, offering on-demand resource allocation and service customisation [393]. However, the inherent resource dynamics and service uncertainties in DSIN render traditional scheduling methods unsuitable for decision-making tasks requiring high efficiency and rapid response, such as service function chain deployment and mapping. Reinforcement learning (RL), as a self-learning and adaptive decision-making approach, has been explored in SDN/NFV networks to some extent [394]. Nevertheless, how to apply RL to integrate lightweight SDN/NFV for enabling adaptive and scalable NRV in DSIN remains relatively scarce, highlighting a gap in the current research.
@@ 5.2 Distributed AI Inference and On-Orbit Information Processing
Given the constrained transmission capacity of satellite-to-ground links, traditional architectures reliant on GS struggle to address the growing demands for massive data return [395, 396]. Moreover, the limited resources of a single satellite render it impractical to process all received data independently. Consequently, leveraging on-orbit computing and storage resources across DSIN becomes essential for the real-time processing of vast sensor data collections. As outlined in Section 2.2, the distributed satellite computing network architecture integrates the available computing, communication, sensing, and storage resources of distributed satellites into a virtual resource pool for unified management. Onboard intelligent fusion of multi-source satellites is achieved by managing multi-dimensional and heterogeneous resources, enabling collaborative distributed computing power. Distributed AI technologies, such as federated learning [397], transfer learning and split learning [398], can be effectively employed within CCS system in DSIN to facilitate on-orbit information processing. Moreover, With the rapid development of AI, large language models (LLMs) have become a star technology in the field of natural language processing. However, the inference process of LLMs often faces significant challenges in terms of computational resources and time costs, especially for on-orbit intelligent applications in satellite networks. To address this issue, distributed AI inference technology has emerged. Distributed AI inference involves breaking down the large language model inference task, which would typically be executed on a single machine, into multiple sub-tasks that can be processed in parallel across multiple nodes with limited computational power. This approach enables the efficient utilisation of computational resources from distributed and edge nodes, significantly improving on-orbit information processing and inference speed, while also reducing inference costs. During the distributed inference offloading process, managing a large number of computational resources, including CPUs, GPUs, and others, is crucial. To accelerate the model inference speed, NRV and distributed satellite computing network architecture, as mentioned in Section $2.2$ and $5.1$, can be employed to achieve compact and efficient computational power collaboration and task priority scheduling. Furthermore, in response to the dynamic network characteristics of DSIN, such as frequent topology changes and narrow communication time windows between the satellites and GSs, distributed incremental inference emerges as a promising technique. By learning the implicit relationships between various tasks or model information, it abstracts a distributed inference model with basic inference capabilities. This approach avoids full-scale learning and model updates, thereby accelerating model inference speed while reducing communication overhead.
@@ 5.3 Semantic Communications in DSIN
The 6G-enabled DSIN are expected to provide ubiquitous intelligent services with stringent QoS requirements in dynamic network environments, which prompts a shift from conventional architectures that focus on high transmission rates between transceivers to new architectures centered on intelligent connectivity of everything. The resulting extensive scales of data which require a massive amount of bits to represent, impose a huge burden and bottleneck on communication systems that handle only bit-level reconstruction, especially on DSIN with limited transmission power, storage space and computing resources [409]. In response, a novel paradigm known as semantic communication is inspired as a potential technology to break the bottleneck by utilizing the semantics and content of data to optimize the utilization of communication resources [410, 411]. For example, when CCS system collaboratively executes a remote sensing task, each satellite extracts semantic information about the target in orbit and transmits it to the ground station or device. This effectively reduces the volume of data transmitted over the link and improves transmission timeliness. In addition, in poor channel conditions, semantic communication proves more robust to fading than traditional bit-level reliable transmission methods, offering greater resilience for DSIN [412, 413]. In this sense, semantic communication can be viewed as a promising development direction for the future of DSIN.
@@ 6 Conclusion
This work was supported in part by the Major Key Project of PCL under Grant PCL2024A01, in part by the National Natural Sciences Foundation of China (NSFC) under Grant 62071141 and Grant 62027802, in part by the Shenzhen Science and Technology Program under Grant JCYJ20241202123904007, Grant GXWD20231127123203001 and Grant JSGG20220831110801003, in part by the Fundamental Research Funds for the Central Universities under Grant HIT.OCEF.2024046, in part by the National Key R&D Program of China No. 2020YFB1806900, and in part by the Beijing Natural Science Foundation No. L242012.
@@
[…]
@@
GRAND Guessing Random Additive SDN Software-Defined Networking
@@
LLMs Large Language Models LPS Low-Correlation-Zone Periodic Sequence
RIS Reconfigurable Intelligent Surfaces